Grade 12th passPhysical Chemistry

1L of N/2 solution was heated in a beaker. When the volume was reduced to 600ml, 9.125g of HCl was lost out. The new normality of solution is approximately?A 0.2B 0.4C 0.8D 0.9

Ashmeet Lamba
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
From the N1V1 = constant concept we get,
1*1/2 = 0.6*N2
we get,
N2 = 10/12 => 5/6 => 0.83
This would be the final normality of the solution.